# Privacy risk in GeoData: A survey

### v1

@@ 4.5. Synthetic data generation
As we described in Section 3.1, synthetic datasets aims to preserve the spatial characteristics and patterns present in the original data while preserving the individuals’ privacy. Quick and Waller (Quick and Waller, 2018) proposes the use of spatiotemporal data analysis methods for generating synthetic data to address disclosure risks associated with releasing public-use data in disease mapping, where small counts present challenges. The study employs a Bayesian model where the synthetic data are generated from the resulting posterior predictive distribution, preserving different dependencies. (Drechsler and Hu, 2021) investigates the viability of using synthetic data generation to provide detailed geocoding information from a large administrative database while maintaining confidentiality. The authors showed that a synthesis strategy based on categorical CART models performs well addressing the risk-utility trade-off.
(Quick, 2021) focuses on the dissemination of synthetic data as an effective means of making sensitive information publicly available while minimising the risk of disclosure. It extends a differential privacy approach to Poisson-distributed count data, accommodating population size heterogeneity and incorporating prior information on event rates. Hong et al. (Hong et al., 2022) proposed a privacy-preserving synthetic data generation framework for addressing privacy concerns associated with sharing complete datasets. The method involves synthesising a new privacy-preserving dataset through a generative model that captures the probability density function of the original dataset’s features. The synthetic dataset is designed to be statistically distinct from the original while preserving data utility.
Koebe et al. (Koebe et al., 2023) proposed an innovative microdata dissemination strategy for household survey data to balance privacy and data utility. Instead of traditional approaches involving perturbed cluster locations, they suggest publishing two datasets: the first contains original survey data with limited geographic identifiers, suitable for representative-level analysis; the second involves synthetically generated survey data with accurate cluster locations, enhancing proximity-related information for spatial analysis. The proposed strategy significantly reduces re-identification risks compared to common spatial perturbation methods, while maintaining data utility for both non-spatial and spatial analyses. The authors employ copulas for synthetic data generation, emphasising their computational efficiency, interpretability, and practicality for data producers. The methodology is seen as beneficial for expanding the use cases of survey data products, supporting mapping initiatives, and ensuring data access while complying with privacy regulations.
@@ 5. Open issues and future directions
As we have reviewed in Section 4, no geomaksing technique is complete and capable of providing high accuracy while ensuring strong privacy against different attacks. Each masking technique has different advantages and disadvantages which requires the data custodians to understand these masking techniques and their capabilities before using these masking techniques in practical applications. While most masking literature provides a detailed explanation of how to use the masking procedure in general, most masking methods lack a corresponding guide for the optimal choice of parameters as these parameter settings need to be changed according to the data that needs protecting. Further, the some papers do not take into account different scenarios that could occur when their proposed masking techniques applied in real applications. For example, a technique such as location swapping could lead into issues when no residential addresses are found within the radius that defined based on the population density of a region (Fabrègue and Bogoni, 2023).
@@
The main question in geomasking research is how to balance the need for anonymisation with the usefulness of geodata, while also considering the probability and potential harms of re-identification (Helderop et al., 2023). In general, geomasking methods could preserve accuracy are presumed to impose a high risk of re-identification. As we discussed in Section 3.1, different approaches to geomsaking, the complete anonymisation is impossible if we want the data to remain useful, and that we should focus research efforts not only on better preventing identification given an attempt, but to also consider the probability of an event and the harms of such incidents.
@@
False identification is an emerging concern of applying geomasking methods, which indicate the linking of the masked data points to incorrect persons or households (Seidl et al., 2018). The false identification transferred the potential negative effects of being identified from the true persons or households to individuals who were not part of the research. Such limitations influence both the disclosure risk and a successful investigation of spatial patterns. There are newly developed geo-masking methods that target addressing this issue, such as adaptive Voronoi masking (Kounadi and Leitner, 2016). Though these techniques might be better approaches to visualising a protected version of the distribution of a point pattern, however, they are less accurate in detecting local patterns in location data. Further, these techniques might not be applicable to different geo data types that are complex, such as spatio-temporal stamps of a user in a social network.
@@
As we have discussed in Section 3.1, many geomasking techniques could not provide provable privacy guarantees as their underlying methodologies only contain non-provable privacy techniques, such as k-anonymity (Wang et al., 2022; Broen et al., 2021). Techniques such as differential privacy has been used in location based services (Yan et al., 2022) for providing strong and provable privacy guarantees when answering queries, but has not yet been adopted or utilised in clustering and aggregation of geodata (Harris, 2020). For example, to improve geodata anonymisation and align with the relevant risk factors, it may have been more appropriate to add random noise to the geographic information of routes mapped by just a few individuals, or exclude routes mapped by very few individuals (e.g., publish the number of routes in the area, but not the exact paths).
@@
In the geomasking literature, each paper tends to use different data sets and different measures to evaluate their proposed methodologies. As shown by Redlich (Redlich, 2022), the choise of different accuracy and privacy metrics influence the evaluation of the geomasking methods leading to different conclusions. This utilisation of different experimental settings in geomasking literature could potentially lead the comparison of different geomasking techniques difficult. Thus, having a benchmark setting, including a set of data sets, accuracy measures, and privacy measures, would allow more accurate quantification of performance of different geomasking methods.
@@
IBM Geospatial foundation model (Raghavan and Shim, 2023)
In collaboration with NASA, IBM has developed the first-ever foundation model for analysing geospatial data, watsonx.ai model, to convert satellite data into high-resolution maps. The model aims to assist in estimating climate-related risks, monitoring forests for carbon-offset programs, and creating predictive models for climate change adaptation. Unlike traditional LLMs requiring separate training for each task, this model, based on NASA’s Harmonized Landsat Sentinel-2 data, is trained to understand satellite images, recognising features like historic floods and fire burn scars.
ChatGeoPT (Strong, 2023)
ChatGeoPT, a Geospatial AI Assistant, serves as a demonstrative prototype that explores the integration of LLMs into geospatial data science workflows. The application accepts natural language prompts and translates them into queries for the OpenStreetMap (OSM). ChatGeoPT is seen as a preliminary step toward the development of a more extensive Large Earth Observation Model (LEOM), aiming to enable intelligent interactions and summarisations of geospatial data from diverse sources, including satellite imagery and weather data, extending beyond the capabilities of OpenStreetMap.
MapsGPT (Proxi, nd)
MapsGPT simplifies the process of creating interactive maps, eliminating the need for extensive coding skills and making it accessible to everyone. Unlike traditional GIS, MapsGPT allows users to generate customised maps effortlessly by inputting a simple sentence. MapsGPT provides a versatile solution for discovering and curating locations, making it easy to plan activities, find essential stores, and explore a new city.
LLM-Geo (Li and Ning, 2023)
LLM-Geo represents an innovative paradigm known as Autonomous GIS, introducing an AI-powered geographic information system that leverages LLMs like ChatGPT to transform spatial analysis. LLM-Geo seeks to automate spatial data collection, analysis, and visualisation by harnessing the natural language understanding, reasoning, and coding capabilities of LLMs.
GEoLM (Li et al., 2023)
GEoLM, a geospatially grounded language model, enhances natural language understanding by integrating valuable geospatial information from extensive databases like OpenStreetMap (OSM). GEoLM connects linguistic details in textual corpora with geospatial data retrieved from geographical databases. GEOLM stands as a versatile tool for various geographically related language understanding tasks, with future plans to extend its applications to question-answering and auto-regressive tasks.
GFM (Mendieta et al., 2023)
GFM developes geospatial foundation models for applications like agriculture, urban planning, and disaster response. The authors introduce GeoPile, a compact and diverse dataset sourced from various remote sensing platforms, to enhance feature diversity during pretraining. Departing from traditional approaches, they explore the potential of continual pretraining from readily available ImageNet-22k models.
@@ 5.1. Integration of large language models (LLMs) with Geospatial data
Rao et al.(Rao et al., 2023) examined the privacy and security risks associated with the entire lifecycle of GeoAI foundation models. The study identifies a range of potential privacy and security risks in different phases of GeoAI foundation model development. The authors highlighted the challenges in using such foundations models in geospatial research mainly include balancing privacy and utility, mitigating geographic biases, and ensuring cross-modal privacy risks are addressed.
Several other studies have explored the use of foundation models in geospatial data analysis. mooney et al. (Mooney et al., 2023) investigated the performance of ChatGPT in a real Geographic Information Systems (GIS) exam. The study identifies challenges in teaching spatial concepts to LLMs, mainly due to the limitations in handling multimodal content. (Tao and Xu, 2023) explore the potential of leveraging ChatGPT for the task of map design. The authors conducted two sets of experiments, testing the generation of thematic maps through specified prompts and the creation of mental maps purely from textual descriptions of geographic spaces. Though ChatGPT’s reliance on textual input and output, the authors showed that ChatGPT holds the potential to revolutionise map design processes, particularly in educational and research contexts, while emphasising the need for cautious consideration of its current limitations and validation of data sources.
Similarly, Zhang et al. (Zhang et al., 2023b) proposed GeoGPT, a framework that integrates the semantic understanding capabilities of LLMs. GeoGPT utilises the Langchain framework to connect the LLM with a pool of GIS tools, allowing it to choose and execute appropriate tools for tasks like geospatial data collection, spatial analysis, and mapping. (Zhang et al., 2023a) explore the ethical dimensions of leveraging generative AI in the domain of cartography. Focusing specifically on maps generated by LLMs, the researchers establish an open-source dataset containing both AI-generated and human-designed maps. The authors showed the LLM generated maps lack with geographic coverage and diversity, and but highlighted that such AI tools could be used in research where real geodata is limited. Table 6 highlights some other tools specifically designed to integrate LLMs with geospatial data.
@@ 6. Conclusion
In this paper we have presented a survey of historical and current state-of-the-art masking techniques for geodata. We have identified 11 dimensions that allowed us to characterise geomasking techniques, and to generate a taxonomy of such techniques. This proposed taxonomy can be used as a comparison and analysis tool for geomasking techniques. Through this taxonomy we identified various shortcomings of current approaches to geomasking that suggest several future research directions in this field.
Crucially, there is currently no overarching framework available that allows different geomasking approaches to be evaluated comparatively. Researchers have used various datasets as well as a variety of measures to evaluate the privacy, scalability, and accuracy, of their geomasking techniques. Most research in geomasking so far has concentrated on the development of anonymisation of geodata, while only limited work has been conducted on exploring how these masking techniques affect false identifications, the use of provable privacy techniques such as differential privacy for geodata clustering and aggregation, and masking techniques that can provide both high accuracy while scalable to large datasets and provide strong privacy. Furthermore, an area for future research involves investigating how large language models (LLMs) can be employed to analyse and enhance geomasking techniques. To date, no studies have delved into the potential integration of LLMs for this purpose, making it a promising area for future exploration and improvement. Solving these open research questions is a core requirement to make geomasking applicable for practical applications.

### v2

@@ 4.5. Synthetic data generation
Quick and Waller (Quick and Waller, 2018) propose the use of spatiotemporal data analysis methods for generating synthetic data to address disclosure risks associated with releasing public-use data in disease mapping, where small counts present challenges. The study employs a Bayesian model where the synthetic data are generated from the resulting posterior predictive distribution, preserving different dependencies. Drechsler and Hu (Drechsler and Hu, 2021) investigate the viability of using synthetic data generation to provide detailed geocoding information from a large administrative database while maintaining confidentiality. The authors show that a synthesis strategy based on categorical CART models performs well in addressing the risk-utility trade-off.
Quick (Quick, 2021) focuses on the dissemination of synthetic data as an effective means of making sensitive information publicly available while minimising the risk of disclosure. It extends a differential privacy approach to Poisson-distributed count data, accommodating population size heterogeneity and incorporating prior information on event rates. Hong et al. (Hong et al., 2022) propose a privacy-preserving synthetic data generation framework for addressing privacy concerns associated with sharing complete datasets. The method involves synthesising a new privacy-preserving dataset through a generative model that captures the probability density function of the original dataset’s features. The synthetic dataset is designed to be statistically distinct from the original while preserving data utility.
Koebe et al. (Koebe et al., 2023) propose an innovative microdata dissemination strategy for household survey data to balance privacy and data utility. Instead of traditional approaches involving perturbed cluster locations, they suggest publishing two datasets: the first contains original survey data with limited geographic identifiers, suitable for representative-level analysis; the second involves synthetically generated survey data with accurate cluster locations, enhancing proximity-related information for spatial analysis. The proposed strategy significantly reduces re-identification risks compared to common spatial perturbation methods, while maintaining data utility for both non-spatial and spatial analyses. The authors employ copulas for synthetic data generation, emphasising their computational efficiency, interpretability, and practicality for data producers. The methodology is seen as beneficial for expanding the use cases of survey data products, supporting mapping initiatives, and ensuring data access while complying with privacy regulations.
@@ 5. Open issues and future directions
As we have reviewed in Section 4, no geomasking technique is complete and capable of providing high accuracy while ensuring strong privacy against different attacks. Each masking technique has distinct advantages and disadvantages, requiring data custodians to understand these techniques and their capabilities before applying them in practical applications. While most masking literature provides a detailed explanation of how to use the masking procedure in general, most masking methods lack a corresponding guide for the optimal choice of parameters, as these parameter settings need to be changed according to the data that needs protection. Further, some papers do not take into account different scenarios that could occur when their proposed masking techniques are applied in real applications. For example, a technique such as location swapping could lead to issues when no residential addresses are found within the radius that is defined based on the population density of a region (Fabrègue and Bogoni, 2023).
@@
The main question in geomasking research is how to balance the need for anonymisation with the usefulness of geodata while also considering the probability and potential harms of re-identification (Helderop et al., 2023). In general, geomasking methods can preserve data accuracy but are presumed to impose a high risk of re-identification. As discussed in Section 3.1, complete anonymisation is impossible if the data is to remain useful. Therefore, research efforts should focus not only on better preventing identification but also on considering the probability of re-identification events and the potential harms of such incidents.
@@
False identification is an emerging concern of applying geomasking methods, which indicate the linking of the masked data points to incorrect persons or households (Seidl et al., 2018). The false identification transferred the potential negative effects of being identified from the true persons or households to individuals who were not part of the research. Such limitations influence both the disclosure risk and a successful investigation of spatial patterns. There are newly developed geomasking methods that target addressing this issue, such as adaptive Voronoi masking (Kounadi and Leitner, 2016). Though these techniques might be better approaches to visualising a protected version of the distribution of a point pattern, however, they are less accurate in detecting local patterns in location data. Further, these techniques might not be applicable to different geodata types that are complex, such as spatiotemporal stamps of a user in a social network.
@@
As we have discussed in Section 3.1, many geomasking techniques could not provide provable privacy guarantees as their underlying methodologies only contain non-provable privacy techniques, such as k-anonymity (Wang et al., 2022; Broen et al., 2021). Techniques such as differential privacy have been used in location-based services (Yan et al., 2022) for providing strong and provable privacy guarantees when answering queries, but have limited adoption and utilisation in the masking of geodata (Harris, 2020). For example, to improve geodata anonymisation and align with the relevant risk factors, it may have been more appropriate to add random noise to the geographic information of routes mapped by just a few individuals, or exclude routes mapped by very few individuals (e.g., publish the number of routes in the area, but not the exact paths).
@@
In the geomasking literature, each paper tends to use different datasets and different measures to evaluate their proposed methodologies. As shown by Redlich (Redlich, 2022), the choice of different accuracy and privacy metrics influence the evaluation of the geomasking methods, leading to different conclusions. The utilisation of different experimental settings in geomasking literature could potentially make the comparison of different geomasking techniques difficult. Thus, having a benchmark setting, including a set of datasets, accuracy measures, and privacy measures, would allow for more accurate quantification of the performance of different geomasking methods.
@@
IBM Geospatial foundation model  (Raghavan and Shim, 2023) In collaboration with NASA, IBM has developed the first-ever foundation model for analysing geospatial data, watsonx.ai model, to convert satellite data into high-resolution maps. The model aims to assist in estimating climate-related risks, monitoring forests for carbon-offset programs, and creating predictive models for climate change adaptation. This model, based on NASA’s Harmonised Landsat Sentinel-2 data, is trained to understand satellite images, recognising features like historic floods and fire burn scars.
ChatGeoPT (Strong, 2023) ChatGeoPT, a Geospatial AI Assistant, is a prototype that explores the integration of LLMs into geospatial data science workflows. The application accepts natural language prompts and translates them into queries for the OpenStreetMap (OSM). ChatGeoPT is seen as a preliminary step toward the development of a more extensive Large Earth Observation Model (LEOM), aiming to enable intelligent interactions and summarisations of geospatial data from diverse sources, including satellite imagery and weather data, extending beyond the capabilities of OpenStreetMap.
MapsGPT (Proxi, nd) MapsGPT simplifies the process of creating interactive maps, eliminating the need for extensive coding skills and making it accessible to everyone. Unlike traditional GIS, MapsGPT allows users to generate customised maps effortlessly by inputting a simple sentence. MapsGPT provides a versatile solution for discovering and curating locations, making it easy to plan activities, find essential stores, and explore a new city.
LLM-Geo (Li and Ning, 2023) LLM-Geo represents an innovative paradigm known as Autonomous GIS, introducing an AI-powered geographic information system that leverages LLMs like ChatGPT to transform spatial analysis. LLM-Geo seeks to automate spatial data collection, analysis, and visualisation by harnessing the natural language understanding, reasoning, and coding capabilities of LLMs.
GEoLM (Li et al., 2023) GEoLM, a geospatially grounded language model, enhances natural language understanding by integrating valuable geospatial information from extensive databases like OSM. GEoLM connects linguistic details in textual corpora with geospatial data retrieved from geographical databases. GEOLM stands as a versatile tool for various geographically related language understanding tasks, with future plans to extend its applications to question-answering and auto-regressive tasks.
GFM (Mendieta et al., 2023) GFM develops geospatial foundation models for applications like agriculture, urban planning, and disaster response. The authors introduce GeoPile, a compact and diverse dataset sourced from various remote sensing platforms, to enhance feature diversity during pretraining. They explore the potential of continual pretraining from readily available ImageNet-22k models.
@@ 5.1. Integration of large language models (LLMs) with Geospatial data
Rao et al.(Rao et al., 2023) examine the privacy and security risks associated with the entire lifecycle of GeoAI foundation models. The study identifies a range of potential privacy and security risks in different phases of GeoAI foundation model development. The authors highlight that leveraging such foundational models in geospatial research poses challenges, particularly in balancing privacy and utility, mitigating geographic biases, and addressing cross-modal privacy risks.
Several other studies have explored the use of foundation models in geospatial data analysis. Mooney et al. (Mooney et al., 2023) investigate the performance of ChatGPT in a real Geographic Information Systems (GIS) exam. The study identifies challenges in teaching spatial concepts to LLMs, mainly due to the limitations in handling multimodal content. Tao and Xu (Tao and Xu, 2023) explore the potential of leveraging ChatGPT for the task of map design. The authors conduct two sets of experiments, testing the generation of thematic maps through specified prompts and the creation of mental maps purely from textual descriptions of geographic spaces. Though ChatGPT’s reliance on textual input and output, the authors show that ChatGPT holds the potential to revolutionise map design processes, particularly in educational and research contexts, while emphasising the need for cautious consideration of its current limitations and validation of data sources.
Similarly, Zhang et al. (Zhang et al., 2023b) propose GeoGPT, a framework that integrates the semantic understanding capabilities of LLMs. GeoGPT utilises the Langchain framework to connect the LLM with a pool of GIS tools, allowing it to choose and execute appropriate tools for tasks like geospatial data collection, spatial analysis, and mapping. Zhang et al. (Zhang et al., 2023a) explore the ethical dimensions of leveraging generative AI in the domain of cartography. Focusing specifically on maps generated by LLMs, the researchers establish an open-source dataset containing both AI-generated and human-designed maps. The authors show that the maps generated by the LLM lack geographic coverage and diversity. However, they highlight that such AI tools could be valuable in research settings where real geodata is limited. Table 6 highlights some other tools specifically designed to integrate LLMs with geospatial data.
@@ 6. Conclusion
In this paper, we present a survey of historical and current state-of-the-art masking techniques for geodata. We proposed a taxonomy that serves as a comparison and analysis tool for geomasking techniques. In this taxonomy we identified three main dimensions: privacy techniques, evaluation measures, and practical aspects. These dimensions enabled us to characterize geomasking techniques and generate a taxonomy of such techniques. Through this taxonomy, we identified various shortcomings in current approaches to geomasking, suggesting several future research directions in this field. An area for future research involves investigating how large language models (LLMs) can be employed to analyse and enhance geomasking techniques. To date, no studies have delved into the potential integration of LLMs for this purpose, making it a promising area for future exploration and improvement. Solving these open research questions is a core requirement to make geomasking applicable for practical applications.